OpenStudy (carniel):

Brain teaser :) 6 + 4 = 210 9 + 2 = 711 5 + 3 = 208 1st person to get it right gets a medal

OpenStudy (carniel):

Wrong

OpenStudy (anonymous):

.-.

OpenStudy (carniel):

I wouldn't have = that number if it was those...

OpenStudy (anonymous):

6 + 4 = 10 9 + 2 = 11 5 + 3 = 8

OpenStudy (anonymous):

21 64 26

OpenStudy (carniel):

Wrong :O x2

OpenStudy (anonymous):

why?

OpenStudy (carniel):

6 + 4 = 210.... Stop trying to get a different answer at the end. ._.

OpenStudy (anonymous):

Just nvm .-.

OpenStudy (anonymous):

6-4 = 2 6+4 = 10 (210) 9-2 = 7 9+2 = 11 (711) 5-3 = 2 5+3 = 08 (208)

OpenStudy (carniel):

CORRECT

OpenStudy (carniel):

:O Smart person

OpenStudy (anonymous):

Sweet! :D

OpenStudy (carniel):

Everyone please give him a medal ^^

OpenStudy (kewlgeek555):

That moment when you're like "OOOHHHHHHHHHHHH"

OpenStudy (carniel):

LoL

OpenStudy (geserino):

Okay, why 5 + 3 = 208 and not 37 ?

OpenStudy (geserino):

I mean, where did that zero came from?

OpenStudy (anonymous):

6+4=10 (210) The 2 before the 10 (210) That means that 4 is 2 numbers below 6! Test with the rest 7+6=13 (113) See 1 it means 6 is 1 number from 7! Boom nailed it